Troy Engines, Ltd., manufactures a variety of engines for use in heavy equipment. The company has always produced all of theComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Assuming the cComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Should the outComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Suppose that iComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 ---- Given the
Answer #1

1) Differential analysis

Make Buy
Direct material 270000
Direct labor 162000
variable overhead 72000
Fixed overhead (108000/3) 36000
Purchase cost (18000*33) 594000
Total relevant cost 540000 594000

Financial (disadvantage) = 540000-594000 = -54000

2) No

3) Differential analysis

Make Buy
Direct material 270000
Direct labor 162000
variable overhead 72000
Fixed overhead (108000/3) 36000
Opportunity Cost 180000
Purchase cost (18000*33) 594000
Total relevant cost 720000 594000

Financial advantage = 720000-594000 = 126000

4) Yes